Answer the question on the basis of the following demand and cost data for a specific firm.

Demand Data Cost Data
(1) Price (2) Price (3) Quantity Output Total Cost
$11.00 $10.00 6 6 $61
9.99 8.85 7 7 62
9.00 8.00 8 8 64
8.00 7.00 9 9 67
7.10 6.10 10 10 72
6.00 5.00 11 11 79
5.15 4.15 12 12 86

If columns (1) and (3) of the demand data shown are this firm's demand schedule, the profit-maximizing level of output will be

Multiple Choice

  • 12 units.

  • 8 units.

  • 10 units.

  • 9 units.
Answer #1

Answer: 8 units

(1) Price

(3) Quantity

Total revenue ($)

Marginal revenue ($)

Total Cost

Marginal cost ($)

$11.00

6

66

----

$61

----

9.99

7

69.93

3.93

62

1

9.00

8

72

2.07

64

2

8.00

9

72

0

67

3

7.10

10

71

-1

72

5

6.00

11

66

-5

79

7

5.15

12

61.8

-4.2

86

7

Profit maximization condition:

Profit is maximum when (MC) Marginal cost = (MR) Marginal revenue

MC = MR = $2 at 8th unit of output.

Formulas used:

Total revenue = Price * Quantity

Marginal revenue = Total revenue of n – Total revenue of n-1

Marginal cost = Total cost of n – Total cost of n-1

Calculation:

(1) Price

(3) Quantity

Total revenue ($)

Marginal revenue ($)

Total Cost

Marginal cost ($)

$11.00

6

11*6=66

----

$61

----

9.99

7

9.99*7=69.93

69.93-66=3.93

62

62-61=1

9.00

8

9*8=72

72-69.93=2.07

64

64-62=2

8.00

9

8*9=72

72-72=0

67

67-64=3

7.10

10

7.10*10=71

71-72=-1

72

72-67=5

6.00

11

6*11=66

66-71=-5

79

79-72=7

5.15

12

5.15*12=61.8

61.8-66=-4.2

86

86-79=7